# From Imitation to Refinement Residual RL for Precise Visual Assembly

### v1

@@ 6 Limitations and Conclusion
The local nature of our residual policies is not well suited for learning the macro-level corrective behaviors that are required to recover from large-scale deviations like dropped parts and out-of-distribution in-hand object poses. Furthermore, despite showcasing the advantage of incorporating simulation data, we have yet to achieve high absolute success rates in the real world. Finally, our proposed pipeline struggles in regimes with very high initial scene randomness, as both the base policies and actions produced via RL exploration struggle to deal with out-of-support initial part poses. Exciting future investigations may include better sim-to-real transfer techniques, exploration mechanisms for discovering how to correct large-scale execution errors and incorporating inductive biases that help with generalization to much broader initial state distributions.
@@ Conclusion
This work presents an approach for fine-tuning BC-trained policies for precise manipulation tasks using sparse rewards we call ResiP. We use RL to train residual policies that produce locally corrective actions on top of base models with architecture components that complicate RL, such as diffusion models and chunked action predictions. Our results show the proposed method outperforms alternative techniques for fine-tuning imitation-learned assembly policies with RL. We furthermore show that through teacher-student distillation and sim-to-real co-training techniques, the precise behaviors acquired by our residual learner can be distilled into a real-world assembly policy that operates directly from RGB images.
Appendix
@@ A.1 Training Hyperparameters
We provide a detailed set of hyperparameters used for training. General hyperparameters for all models can be found in Tab. 3, while specific hyperparameters for the diffusion models are in Tab. 4, and those for the MLP baseline are in Tab. 5.
Table 3: Training hyperparameters shared for all state-based BC models
@@
Table 4: State-based diffusion pre-training hyperparameters
@@
Table 5: State-based MLP pre-training hyperparameters
@@
Below, we list the hyperparameters used for online reinforcement learning fine-tuning. The parameters that all state-based RL methods methods shared are in Tab. 6. Method-specific hyperparameters for training the different methods are in the tables below, direct fine-tuning of the MLP in Tab. 7, online IDQL in Tab. 8, and the residual policy in Tab. 9. The different methods were tuned independently, but the same hyperparameters were used for all tasks within each method.
Table 6: Hyperparameters shared for all online fine-tuning approaches
@@
Table 7: Hyperparameters for direct fine-tuning of MLP
@@
Table 8: Hyperparameters for training value-augmented diffusion sampling (IDQL)
@@
Table 9: Hyperparameters for residual PPO training
@@
Actor Gaussian initial log st.dev. -1.5
@@
For the real-world experiments, we use a separate set of hyperparameters, presented in Tab. 10. The main difference is that we found in experimentation that the transformer backbone in  [14] worked better than the UNet for real-world experiments. These models are also operating from RGB observations instead of privileged states, and we provide parameters for the image augmentations applied to the front camera in Tab. 11 and the wrist camera in Tab. 12.
Table 10: Training hyperparameters for real-world distilled policies
@@
Table 11: Parameters for front camera image augmentation
@@
Table 12: Parameters for front camera image augmentation
@@
A.2 Action and State-Space Representations
@@ Appendix B Tasks and Environment
B.1 Tasks details and reward signal
We detail a handful of differentiating properties for each of the three tasks we use in Tab. 13. one_leg involves assembling 2 parts, the tabletop and one of the 4 table legs. The assembly is characterized as successful if the relative poses between the parts are close to a predefined assembled relative pose. When this pose is achieved, the environment returns a reward of 1. That is, for the one_leg task, the policy received a reward of 1 only at the very end of the episode. For round_table and lamp, which consists of assembling 3 parts together, the policy receives a reward signal of 1 for each pair of assembled parts. E.g. for the lamp task, when the bulb is fully screwed into the base, the first reward of 1 is received, and the second is received when the shade is correctly placed.
Table 13: Task Attribute Overview
One Leg Round Table Lamp
Mean episode length $\sim$500 $\sim$700 $\sim$600
# Parts to assemble 2 3 3
Num rewards 1 2 2
Dynamic bbject
# Precise insertions 1 2 1
# Screwing sequences 1 2 1
@@ B.1 Tasks details and reward signal
B.2 Details on randomization scheme
The “low” and “medium” randomness settings we used for data collection and evaluation reflect how much the initial part poses may vary when the environment is reset. We tuned these conditions to mimic the levels of randomness introduced in the original FurnitureBench suite [15]. However, we found that their method of directly sampling random poses often leads to initial part configurations that collide with each other, requiring expensive continued sampling to eventually find an initial layout where all parts do not collide with each other.
Our modified randomization scheme instead initializes parts to a single pre-specified set of feasible configurations and then applies a randomly sampled force and torque to each part (where the force/torque magnitudes are tuned for each part and scaled based on the desired level of randomness). This scheme allows the physics simulation to ensure parts stay out of collision while still providing a controlled amount of variation in the initial scene randomness.
The second way we modified the randomization scheme was to randomize the position of the U-shaped obstacle fixture and the parts (the obstacle fixture was always kept in a fixed position in [15]). Our reasoning was that, for visual sim-to-real without known object poses, we could only imperfectly and approximately align the obstacle location in the simulated and real environment. Rather than attempting to make this alignment perfect, we instead trained policies to cover some range of possible obstacle locations, hoping that the real-world obstacle position would fall within the distribution the policies have seen in simulation. Fig. 8 shows examples of our different randomness levels for each task in simulation.
Figure 8: Examples of initial scene layouts for one_leg, lamp, and round_table with different levels of initial part pose and obstacle fixture randomness
B.3 Adjustments to FurnitureBench simulation environments
In addition to our modified force-based method of controlling the initial randomness, we introduced multiple other modifications to the original FurnitureBench environments proposed in [15] to enable the environment to run fast enough to be feasible for online RL training. With these changes, we were able to run at a total $\sim$4000 environment steps per second across 1024 parallel environments. The main changes are listed below:
@@ Appendix E Real-World Results
E.1 Quantitative Results Failure Mode Breakdown
@@
The diagram in Fig. 12 shows how successful and failed completion of individual sub-skills along the one_leg task amount to our overall final success rates reported in Tab. 2 (bottom row, corresponding to “40 real + 350 sim” with random initial part poses and a fixed obstacle pose).
E.2 Extension of Pipeline to Unseen Settings

### v3

@@ VI Limitations and Conclusion
The local nature of our residual policies is not well-suited for learning the macro-level behaviors required to recover from large-scale deviations like dropped parts. Our proposed method also struggles in regimes with very high initial scene randomness, as both the base policies and actions produced via RL exploration struggle to deal with out-of-support initial part poses. Furthermore, despite showcasing the advantage of incorporating simulation data, sim-to-real for RGB policies still presents a challenge. There remains a performance gap in both teacher-student distillation and sim-to-real distribution shifts. Future investigations may include better sim-to-real transfer techniques, exploration mechanisms for discovering how to correct large-scale execution errors, tractable interactive learning for real-world policy distillation, and incorporating inductive biases that help generalize to much broader initial state distributions.
@@ Conclusion
This work presents ResiP, a simple yet effective approach for improving behavior-cloned policies on precise manipulation tasks using sparse rewards. Our key insight is recognizing modern chunked BC policies as trajectory planners that require closed-loop corrections for reliable execution. By training a residual policy with RL to provide these corrections, we sidestep the complexities of directly fine-tuning architectures like diffusion models while addressing fundamental limitations of BC – distribution shift and lack of reactivity.
Our results demonstrate that this approach significantly outperforms pure BC methods and direct RL fine-tuning across multiple challenging assembly tasks, improving success rates from $\sim$50-80to over 95% in many cases. Through careful ablations, we show the importance of the fully closed-loop nature of the residual and the choice of a robust base policy, particularly for high-precision tasks like sub-millimeter insertions.
While we demonstrate successful real-world transfer through teacher-student distillation and sim-to-real co-training, the remaining performance gap suggests opportunities for improved sim-to-real procedures through direct vision-based RL or interactive distillation. Additionally, our results with different base architectures suggest that this residual learning approach is valuable for fine-tuning foundation models while preserving diverse capabilities.
Acknowledgments
This work was partly supported by the Sony Research Award and the US Government. The computations in this paper were run on the FASRC cluster, supported by the FAS Division of Science Research Computing Group at Harvard University, and on the MIT Supercloud[88]. Experiment tracking and model checkpoint storage were provided by Weights and Biases.
Author Contributions
LA led the project and implemented most of the code and training infrastructure, including the main residual PPO implementation, and is the primary author.
AS helped conceive of and advise on project goals, led deployment on real hardware and sim-to-real rendering, and helped write the paper.
IS led the implementation of reinforcement learning baselines, helped debug residual PPO implementation, and helped write the paper.
MT provided valuable insights, recommendations, and discussions on reinforcement learning fine-tuning and sim-to-real transfer.
PA advised the project and provided valuable feedback on project framing and contributions.
@@ A-A Training Hyperparameters
We provide a detailed set of hyperparameters used for training. General hyperparameters for all models can be found in Tab. III, while specific hyperparameters for the diffusion models are in Tab. IV, and those for the MLP baseline are in Tab. V.
TABLE III: Training hyperparameters shared for all state-based BC models
@@
TABLE IV: State-based diffusion pre-training hyperparameters
@@
TABLE V: State-based MLP pre-training hyperparameters
@@
Below, we list the hyperparameters used for online reinforcement learning fine-tuning. The parameters that all state-based RL methods methods shared are in Tab. VI. Method-specific hyperparameters for training the different methods are in the tables below, direct fine-tuning of the MLP in Tab. VII, online IDQL in Tab. VIII, and the residual policy in Tab. IX. The different methods were tuned independently, but the same hyperparameters were used for all tasks within each method.
TABLE VI: Hyperparameters shared for all online fine-tuning approaches
@@
TABLE VII: Hyperparameters for direct fine-tuning of MLP
@@
TABLE VIII: Hyperparameters for training value-augmented diffusion sampling (IDQL)
@@
TABLE IX: Hyperparameters for residual PPO training
@@
Actor Gaussian initial log st.dev. -1.0
@@
For the real-world experiments, we use a separate set of hyperparameters, presented in Tab. X. The main difference is that we found in experimentation that the transformer backbone in  [18] worked better than the UNet for real-world experiments. These models are also operating from RGB observations instead of privileged states, and we provide parameters for the image augmentations applied to the front camera in Tab. XI and the wrist camera in Tab. XII.
TABLE X: Training hyperparameters for real-world distilled policies
@@
TABLE XI: Parameters for front camera image augmentation
@@
TABLE XII: Parameters for wrist camera image augmentation
@@
A-B Action and State-Space Representations
@@ Appendix B Tasks and Environment
B-A Tasks details and reward signal
Furniture assembly tasks
We detail a handful of differentiating properties for each of the three tasks we use in Tab. XIII. one_leg involves assembling 2 parts, the tabletop and one of the 4 table legs. The assembly is characterized as successful if the relative poses between the parts are close to a predefined assembled relative pose. When this pose is achieved, the environment returns a reward of 1. That is, for the one_leg task, the policy received a reward of 1 only at the very end of the episode. For round_table and lamp, which consists of assembling 3 parts together, the policy receives a reward signal of 1 for each pair of assembled parts. E.g. for the lamp task, when the bulb is fully screwed into the base, the first reward of 1 is received, and the second is received when the shade is correctly placed.
Real-to-sim task: mug-rack
This task involves the robot picking up a coffee mug and hanging it by the handle on one of two pegs on a rack. See Fig. 13 for task illustration. This task is interesting for two main reasons. First, we don’t have any CAD models for the objects. Instead, we used scanned imports of real-world objects (obtained with the ARCode app on the iPhone App Store). Second, the task has inherent multi-modality in that the mug can be hung in one of two ways for each of the two pegs.
The diffusion and residual policy system works well for this task. First, the base diffusion model captures the task’s multimodality and sometimes hangs the mug on both pegs. Furthermore, the residual RL procedure keeps this multimodality intact as the base model is frozen.
(a) Example task initialization of the mug-rack task.
(b) Example of hanging the mug on the lower rack.
(c) Example of hanging the mug on the upper rack.
Figure 13: Overview of the mug-rack task to showcase the real-to-sim capabilities one can leverage with our pipeline. This also shows how reward signals can be inferred directly from data instead of being hand-designed. Finally, as the task can be completed in one of several ways, this task also tests the policies’ capability to deal with multi-modality.
High-precision, Factory task: peg-in-hole
To push the limits of precision in simulation, controller, and policy, we pick one of the insertion tasks from the Factory task suite [49], which involves grasping a peg and inserting in a hole with a 0.2mm clearance, i.e., 25x tighter than the FurnitureBench [5] tasks. See Fig. 14 for task illustration.
Our approach also worked out of the box on this task, using the same hyperparameters as for the FurnitureBench tasks. Here, we achieve $5\%$ success rate in pre-training and $\sim$99in fine-tuning. Good performance at this task is essentially entirely dominated by the ability to locally adjust the peg until it lines up with the hole, and the high final success rate achieved by our approach reflects that the local nature of the corrections learned by our residual policy is well aligned with such task scenarios.
(a) Example task initialization of the peg-in-hole task.
(b) Example of task completion when the peg is fully inserted.
Figure 14: Overview of the peg-in-hole task we add to push the requirement for precision. We find that the pipeline as presented works well with the same hyperparameters used for the furniture tasks.
Bimanual, high-precision task: biman-insert
To test whether our method, ResiP, also works for precise tasks with larger action spaces, we create a simple bimanual industrial assembly task that we term biman-insert. See Fig. 15 for an example initial and final state and Fig. 17 for several random initial states. We design the task by creating simple meshes and importing them into the MuJoCo [53] physics engine. We demonstrate the task using the augmented reality-based teleoperation interface DART [50]. All subsequent training uses the same code and hyperparameters as all the other tasks. This task has a relatively short horizon but has a 20-dimensional action space and relatively tight insertion tolerances. We also perform this task at 50 Hz for policy control, showing that our approach is quite general.
(a) Example task initialization of the biman-insert task.
(b) Example of task completion when the plate is fully inserted.
Figure 15: Overview of the biman-insert task we add to push the requirement for precision and bimanual coordination. We find that the pipeline as presented works well with the same hyperparameters used for the furniture tasks and that the increased action space poses no problem for mastering the task.
TABLE XIII: Task Attribute Overview
one_leg round_table lamp mug-rack peg-in-hole biman-insert
Mean episode length $\sim$500 $\sim$700 $\sim$600 $\sim$150 $\sim$200 $\sim$400
# Parts to assemble 2 3 3 2 2 2
Num rewards 1 2 2 1 1 1
Dynamic object
# Precise insertions 1 2 1 0 1 1
# Screwing sequences 1 2 1 0 0 0
@@ B-A Tasks details and reward signal
Control frequency 10 Hz 10 Hz 10 Hz 10 Hz 10 Hz 50 Hz
Degrees-of-Freedom 7 7 7 7 7 14
B-B Details on randomization scheme
The “low” and “medium” randomness settings we used for data collection and evaluation reflect how much the initial part poses may vary when the environment is reset. We tuned these conditions to mimic the levels of randomness introduced in the original FurnitureBench suite [5]. However, we found that their method of directly sampling random poses often leads to initial part configurations colliding, requiring expensive continued sampling to eventually find an initial layout where all parts do not collide.
Our modified randomization scheme instead initializes parts to a single pre-specified set of feasible configurations. Then, it applies a randomly sampled force and torque to each part (where the force/torque magnitudes are tuned for each part and scaled based on the desired level of randomness). This scheme allows the physics simulation to ensure parts stay out of collision while providing a controlled amount of variation in the initial scene randomness.
The second way we modified the randomization scheme was to randomize the position of the U-shaped obstacle fixture and the parts (the obstacle fixture was always kept in a fixed position in [5]). We reasoned that, for visual sim-to-real without known object poses, we could only imperfectly and approximately align the obstacle location in the simulated and real environment. Rather than attempting to make this alignment perfect, we instead trained policies to cover some range of possible obstacle locations, hoping that the real-world obstacle position would fall within the distribution the policies have seen in simulation. Fig. 16 shows examples of our different randomness levels for each task in simulation.
Figure 16: Examples of initial scene layouts for the tasks from the FurnitureBench task suite [5], one_leg, lamp, and round_table, with different levels of initial part pose and obstacle fixture randomness.
Figure 17: Examples of initial scene layouts for the 3 non-FurnitureBench tasks, mug-rack, peg-in-hole, and biman-insert, for their default level of randomness.
B-C Adjustments to FurnitureBench simulation environments
In addition to our modified force-based method of controlling the initial randomness, we introduced multiple other modifications to the original FurnitureBench environments proposed in [5] to enable the environment to run fast enough to be feasible for online RL training. With these changes, we could run at a total of $\sim$4000 environment steps per second across 1024 parallel environments. The main changes are listed below:
@@ Appendix E Real-World Results
E-A Quantitative Results Failure Mode Breakdown
@@
The diagram in Fig. 21 shows how successful and failed completion of individual sub-skills along the one_leg task amount to our overall final success rates reported in Tab. II (bottom row, corresponding to “40 real + 350 sim” with random initial part poses and a fixed obstacle pose).
E-B Extension of Pipeline to Unseen Settings
@@ Appendix F Expanded Related Work
Appendix G Extended Limitations and Further Work
Real-world distillation
Our experiments have demonstrated the effectiveness of online learning versus offline or passive learning through behavior cloning. Still, we employ only offline learning in our teacher-student distillation phase for sim-to-real transfer, which will likely upper-bound the performance we can transfer to the real world. Combining our pipeline with techniques for online learning could improve performance significantly. However, at this point, there are significant challenges to overcome to make this practically applicable to the tasks studied herein.
The field is progressing rapidly, and we are excited to investigate how online learning in the real world can be made practical for a broader set of tasks with longer horizons and less obvious ways of performing automatic state resets in follow-up work. This effort further ties into a more general framework for pre-training and adaptation of robot systems where the deployed robot can continue learning and adapting “on the job” after deployment. These investigations complement the methods presented in this paper and are not in scope.
At the same time, our results indicate that making more capable systems only through increasing the collection of real-world demos may also be fundamentally limited unless online learning is introduced as a fine-tuning step in those systems.
Locality of online correction learning
Though effective, we re-emphasize that our residual online reinforcement learning framework has the fundamental limitation of being bound to the pre-trained policy and mainly performing locally corrective actions. This limitation is both a strength and a weakness. First, the strong pre-trained prior allows RL to perform the tasks and improve, and having a frozen prior helps stabilize training and prevent collapse. At the same time, the degree to which online learning can generalize to states far from the training set is limited.
Limitations of simulators in contact-rich tasks
We have added an experiment for a task from the Factory [49] task suite that pushes the accuracy of the simulator more than with the original FurnitureBench [5] tasks. This new task has a clearance of 0.2mm for the insertion, which shows that the general BC + Residual RL framework also works well in this setting. We did not show, however, that this transfers to the real world, and it would likely be more challenging than in the original tasks for at least two reasons. First, with increased precision requirements, accurate calibration of physics parameters between the actual and simulated environment will likely matter more. Second, performing manipulation from vision when parts are smaller is more challenging.
Appendix H Further Analysis of Offline versus Online Learning
H-A Distillation scaling analysis
The scaling analyses in Fig. 24 show the same trends as in LABEL:fig:overview (right) and Fig. 7 for the tasks lamp and peg-in-hole. We believe that this data point suggests that pure offline learning from demonstrations may not be sufficient for policies to learn robust and reactive policies, pointing towards the necessity for techniques like RL to reach a high level of robustness and reliability.
(a) Scaling analysis for the lamp task. This task appears significantly more conducive to offline learning than the other tasks tested.
(b) Success rates in exploration phase of training for one_leg, low randomness.
Figure 24: We run similar scaling analyses as in LABEL:fig:overview (right) and Fig. 7 for lamp and peg-in-hole. The general findings of interactive learning are that it is more efficient and has higher asymptotical performance. However, the difference appears to be much smaller for lamp and bigger for peg-in-hole. What drives these differences are left for future work.
H-B Interactive distillation with DAgger
DAgger [1] can learn from scratch significantly more efficiently than pure BC measured in both gradient steps and samples. We have added the DAgger performance to the scaling plot, shown in orange in Fig. 24. Consider the scaling plot in LABEL:fig:overview (right) as an example. In $\sim$10k gradient steps, DAgger surpasses BC from 50 human demos trained with $\sim$100k steps. After 10k steps, it has around 800 rollouts in the aggregated dataset. After around 20k gradient steps, it seems to surpass the best-performing BC distillation runs using more than 10k rollouts and 500k gradient steps, at which point it has $\sim$1.5k demonstrations in the replay buffer. This result highlights the effectiveness of online and interactive learning as opposed to learning purely passively from an offline dataset. Furthermore, it highlights that the expert we query is an effective teacher. It also highlights that for interactive learning to be effective, one needs to have a teacher ready to be queried as learning progresses.
Appendix I Residual RL ablations
I-A Effect of fully versus partially closed-loop policies
One differentiating factor of our residual model from some prior work is that the base and residual models make predictions at different frequencies, i.e., every 8 timesteps for the base model and every timestep for the residual model. Making predictions with the most up-to-date information is likely an easier prediction problem, and we expect this to work better than the “standard” setup of letting the residual correct the full output of the base model. When training a residual model that corrects a whole chunk at a time but otherwise uses the same hyperparameters, we observe that training is less sample efficient and performance saturates at a lower success rate. In particular, the chunked residual policy reaches $\sim$85success rate in about 250 million environment steps, while the one-step residual needs about 75 million.
Figure 25: When learning a residual correction term that corrects the whole chunk at a time online, we find that the learning is significantly slower (measured in environment steps) and saturates a lower asymptotic level.
To further probe the difference between fully closed-loop policies and those using chunking, we evaluate the policies with perturbations added to the parts in the environment throughout the episode. In particular, at each timestep, 1% of parts across the environments will have a random force applied to them. The forces are sampled from the same distribution as the initial part randomization distribution.
See Fig. 4 and Tab. XIV for results. We generally see that the partially open-loop policies have a bigger drop in performance when perturbations are introduced, around 20 percentage points compared to 12 for the one-step residual model.
Model No Perturb W/ Perturb Drop in SR
Standard RPPO 98% 86% 12 pp
Chunked RPPO 92% 73% 19 pp
Chunked pre-trained BC 52% 32% 20 pp
DAgger chunked student DP 90% 68% 24 pp
TABLE XIV: Success rates with/without perturbations for different models. SR = Success Rate, pp = percentage points.
I-B Residual base policy ablation
MLP as base
To further tease apart what part of the diffusion policy that provides the most important performance increase, the action chunking or the denoising diffusion process, we run the same residual PPO run for the one_leg task as before, but with the best-performing BC MLP model in place of the diffusion policy. The results are shown in 26(a) and 26(b). The resulting training dynamics are intriguing. Despite the initial success rate of the base model being close to that of the diffusion model, the success rate drops markedly when exploration noise is introduced. This is especially visible in the training performance in plot 2 below. We also notice that the evaluation performance drops as the residual model explores and learns more. However, the residual is eventually able to find actions that the MLP responds better to and, in the end, converges to a similar performance as the diffusion-based runs. In the more challenging task with higher initial state randomness, the same initial dynamic plays out, but the training performance drops to zero, causing the learning to collapse. We conclude that any base model achieving a high enough initial success rate can be plugged into our framework (and, based on our BC experiments, a base model with chunking is likely to outperform one without chunking) but that the expressivity and robustness to input noise offered by diffusion de-noising also contributes to downstream performance benefits during residual RL.
ACT as base
To see if the robustness to noise and suitedness for residual learning is unique to the diffusion-type model, we also implement and test using Action-Chunked Transformer (ACT) [6] as the base model for the one_leg task at low randomness. With some tuning, we find that the ACT model can achieve comparable performance as the diffusion model, though slightly lower, in pre-training. In the fine-tuning phase, however, it functions as well and stably as the diffusion model base, as shown in 26(c). This suggests that the residual RL framework is suitable for a wide range of fine-tuning applications and may be applied to fine-tune even larger and possibly multi-task models.
(a) Evaluation success rates for RL training for one_leg, low randomness for Diffusion and MLP base policy.
(b) Evaluation success rates for RL training for one_leg, medium randomness for Diffusion and MLP base policy.
(c) Evaluation success rates for RL training for one_leg, low randomness for Diffusion and ACT base policy.
Figure 26: We compare the diffusion-based residual RL training performance with the best-performing MLP as the base model in (a) and (b). As we can see, despite having similar pertaining performance (for low randomness), the MLP-based residual model performs poorly compared to the diffusion-based one. On a higher randomness setting, it fails to complete the task. We compare with using the ACT [6] as the base policy in (c). The pre-training performance is slightly worse, but performance quickly catches up during online learning, indicating that the ACT model is also well-suited for residual learning.
I-C Residual action scaling parameter ablation
A design choice we make is the parameter $\alpha=0.1$. The parameter choice is somewhat arbitrary and was informed by some intuitions about the task. For example, since the residual model intends to make local corrections, we want to imbue it with that inductive bias. In the normalized action space, the workspace is constrained to [-1, 1], and letting a $\sigma=1$ for the residual Gaussian model correspond to [-0.1, 0.1] on the macro scale seemed reasonable.
We have tested more values of the parameter $\alpha\in\{0.01,0.05,0.2,1.0\}$, but kept the resulting exploration noise on the macro scale fixed (i.e., scaled with the value of $\alpha$, so $\alpha_{1}\sigma_{1}=\alpha_{2}\sigma_{2}$). The result, shown in the figure below, shows a remarkable robustness to this parameter, and all cases have very similar performance.
We note a couple of observations. First, $\alpha=0.2$ seems to perform slightly better than our original $\alpha=0.1$. Second, different levels of $\alpha$ also result in very different magnitudes of activations at the last layer, which impacts losses. This experiment shows that the resulting performance does not differ significantly, but we suspect it could make training less stable in harder settings.
Figure 27: We test different values of the residual action scaling parameter $\alpha$ and test it for values $\alpha\in\{0.01,0.1,0.2\}$ while adjusting the exploration noise to be such that the macro-level exploration is the same initially. We find that for success rates in this task, the value is not crucial but does cause training dynamics to change, particularly the residual model output norms and policy loss magnitude.